OpenStudy (anonymous):

From a shipment of 500 memory cards, 5 are selected random for inspection. there are 25 defective memory cards in the shipment. a) What is the probability that the sample will contain at least one defective memory card? (write down the formula, no computation yet) B) verify that the binomial distribution can be used to estimate the probability in part (a). (check the that the condition holds) c) use the binomial distribution to estimate the probability that the sample will contain at least one defective memory card.

OpenStudy (kropot72):

a)\[\large P(1\ or\ more\ defective)=1-P(0\ defective)=1-0.95^{5}\]

OpenStudy (anonymous):

.2262

OpenStudy (anonymous):

b and c remain

OpenStudy (kropot72):

b) A variable which has a binomial distribution has the following properties: There is a fixed number of trials. Each trial is independent of the others. Each trial has a constant probability of success. There are only two possible outcomes, 'success' and 'failure' c) \[\large P(1\ or\ more\ defective)=[1-(5C0\times0.05^{0}\times0.95^{5})]\]